A 1740-kg communications satellite is released from a space shuttle to initially orbit the Earth at a radius of 9. 106 m. After being deployed, the satellite's rockets are fired to put it into a higher altitude orbit of radius 4. 107 m. What is the minimum mechanical energy supplied by the rockets to effect this change in orbit?
Answer #1

here initial energy U1 = -G*M*m/r1 = -(6.67*10^-11*5.98*10^24*1740)/(9*10^6) = -7.711*10^10 J...

U2 = -G*M*m/r2 = (6.67*10^-11*5.98*10^24*1740)/(40*10^6) = -1.74*10^10 J...

required energy = U1-U2 = (-7.711*10^10) + (1.74*10^10) = -5.97*10^10 J